OpenStudy (anonymous):

on the first four tests of the term your average is 84%. You think you can score 96% on each of the remaining tests. How many consecutive test scores of 96% would need to bring your average up to 90% for the term? I am learning about rational equations if that helps.

OpenStudy (anonymous):

4

OpenStudy (anonymous):

how

OpenStudy (anonymous):

well the way I did it was the average of 96 and 84 is 90 because (84+96)/2=90 So if you have 4 84's then you have to balance it out with 4 96's. You take 84*4+96*4 and that equals 720 then divide that by the total number of tests which is 8 and that equals 90.

OpenStudy (dumbcow):

\[\frac{84*4 + 96n}{4+n} = 90\] solve for n multiply (4+n) on both sides then distribute and combine like terms \[84*4 + 96n = 90n + 90*4\] \[6n = 4(90-84)\] \[n=4\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Thanks dumbcow I couldn't remember the equation I just did it by logic.

OpenStudy (dumbcow):

no prob